Suppose we wanted to test the hypothesis that a control group of cancer patients (Group 1) would report higher mean pain ratings than an experimental group receiving special massage treatments (Group 2). Use the following information. Compute a t-statistic for independent groups:

mean group 1 = 78.1               SD21 = 42.1     n1 = 25

mean group 2 = 74.1               SD22 = 39.7     n2 = 25

What are the degrees of freedom and the value of t ? Using α=0.05 for a two-tailed test, is this t statistic significant? Show your calculations (or StatCrunch output) for full credit.
Answer #1

Degree of freedom would be = 25+25-2 = 48

Test Statistic

t=(\frac{78.1-74.1}{\sqrt{\frac{42.1^2}{25}+\frac{39.7^2}{25}}})

t = 0.3456

P value = P(t<-0.3456)+P(t>0.3456)

= 0.7311

Since test is greater than 0.05, we failed to reject the null hypothesis